OpenStudy (anonymous):

Nerd is in need,again!!! ima cry if no one replies find all solutions of the equation in the interval [0,2pi) sin(x+3pi/2)-sin(x-3pi/2)=0

OpenStudy (dumbcow):

use sum and difference formulas sin(x+y) = sinxcosy + sinycosx sin(x-y) = sinxcosy - sinycosx

OpenStudy (anonymous):

okay so how would i go about this?

OpenStudy (dumbcow):

well, in this case y=3pi/2 sin 3pi/2 = -1 cos 3pi/2 = 0

OpenStudy (dumbcow):

equation reduces to -cos x -cos x = 0 -2cos x = 0 cos x = 0 x = pi/2, 3pi/2

OpenStudy (anonymous):

thank u dumbcow! :]